OpenStudy (anonymous):

Evaluate the limit by first recognizing the sum as a Riemann sum for a function defined on [0, 1] for lim n-> infinity sigma i=1 7i^3/2n^4

OpenStudy (anonymous):

OKay Nvm I figured it out. Rewrite this as lim(n→∞) Σ(i = 1 to n) (7/2)(i/n)^3 * (1/n) = ∫(0 to 1) (7/2)x^3 dx, with Δx = 1/n and f(x) = (7/2)x^3 (i/n) since (a+(b-a)i/n) = (7/8)x^4 {for x = 0 to 1} USE ANTI-Derivative = 7/8.

OpenStudy (anonymous):

\[\int_0^1 \frac{7 x^3}{2} \, dx=\frac 7 8 \]